OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 8 inches and its perimeter is 46 inches, find the length of the rectangle in inches

OpenStudy (anonymous):

subtract 16 from 46 then divide by two

OpenStudy (lgbasallote):

\[length + length + width + width = perimeter\] \[l + l + 8 + 8 = 46\] \[2l + 16 = 46\] \[2l = 46 - 16\] \[2l = 30\] \[l = ?\]